OpenStudy (anonymous):

find the number of 2*2 matrix satisfying... i) aij is 1 or -1 ii) (a11)2 +(a12)2 = (a21)2 + (a22)2 =2 iii) a11a21 + a12a22 =0

OpenStudy (anonymous):

sounds like a symmetric matrix.

OpenStudy (anonymous):

umm...yea...symmitric or skew symmitric

OpenStudy (anonymous):

i think that's it. Not sure though/

OpenStudy (anonymous):

yea...i think 16 such matrices cn be there as each place cn hv 2 occupantz...hence it shud be 2*2*2*2

OpenStudy (anonymous):

itz a mayb..

OpenStudy (anonymous):

Questions: Why are you doing these puzzle problems? Is it for HW?

OpenStudy (anonymous):

assignment....self practice...

OpenStudy (anonymous):

Doesn't seem useful in my opinion....Idk it's something I know my prof will never ask me for. The only puzzle problem I was asked in my class was one where I had to use cramers rule.

OpenStudy (anonymous):

ver r u frm??

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

umm..no romero

OpenStudy (anonymous):

vese me amritsar se hu n u?/

OpenStudy (anonymous):

chandigarh

OpenStudy (anonymous):

California

OpenStudy (anonymous):

o ohkeyy..m frn india...nd i thnk my teachrz require every typ of such topicz...:(

OpenStudy (anonymous):

Every teaches differently. It's understandable. I know my prof focuses more on theorems and proofs than anything else.

OpenStudy (anonymous):

mmm...yeahh...teachrz put alot in pressure here....sooo..any guesses fr solution??

OpenStudy (anonymous):

2(a11)+(a12)=2 .../2 a11+a12=1 a11=1-a12 a11=1 Or -1from the question. and same for a12. but . 1=1-a12 only when a12=0 ,and -1=1-a12 when a12=2 right . (didn't match the 1st condition) so i don't think there is really any matrix that satisfy those conditions you mentioned. still maybe i am wrong idk.

OpenStudy (anonymous):

umm sry..i typed wrong..i meant (a11)^2 +(a12)^2 = (a21)^2 + (a22)^2 =2.. ..they were square!!